# When can isotropy help adapt LLMs’ next word prediction to numerical domains?

### v3

@@ Abstract
Recent studies have shown that vector representations of contextual embeddings learned by pre-trained large language models (LLMs) are effective in various downstream tasks in numerical domains. Despite their significant benefits, the tendency of LLMs to hallucinate in such domains can have severe consequences in applications such as energy, nature, finance, healthcare, retail and transportation, among others. To guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black-box and provide performance guarantees through explanation. However, there is little theoretical understanding of when pre-trained language models help solve numeric downstream tasks. This paper seeks to bridge this gap by understanding when the next-word prediction capability of LLMs can be adapted to numerical domains through a novel analysis based on the concept of isotropy in the contextual embedding space. Specifically, we consider a log-linear model for LLMs in which numeric data can be predicted from its context through a network with softmax in the output layer of LLMs (i.e., language model head in self-attention). We demonstrate that, in order to achieve state-of-the-art performance in numerical domains, the hidden representations of the LLM embeddings must possess a structure that accounts for the shift-invariance of the softmax function. By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem and providing a performance guarantee. Experiments show that different characteristics of numeric data and model architecture could have different impacts on isotropy.
@@ 1 Introduction
Large language models have demonstrated impressive success in adapting to various downstream tasks in numerical domains, such as finance Garza and Mergenthaler-Canseco (2023); Yu et al. (2023), energy Gao et al. (2024), climate science Jin et al. (2024), healthcare Wang and Zhang (2024), wireless communications Xu et al. (2024), synthetic tabular generation Dinh et al. (2022); Borisov et al. (2023); Xu et al. (2024), among others. Inspired by the success of pre-trained LLMs, several methods have been developed recently in Gruver et al. (2024); Dooley et al. (2023); Nie et al. (2023); Rasul et al. (2024); Woo et al. (2024); Jin et al. (2024); Ansari et al. (2024) by adapting LLM to numerical domains. For many of these numeric downstream tasks, training a linear classifier on top of the hidden-layer representations generated by the pre-trained LLMs have already shown near state-of-the-art performance Jin et al. (2024); Ansari et al. (2024). However, these models in Gruver et al. (2024); Dooley et al. (2023); Nie et al. (2023); Rasul et al. (2024); Woo et al. (2024); Jin et al. (2024); Ansari et al. (2024) are treated as ‘black-box’ where numeric forecasts are controlled by complex nonlinear interactions between many parameters. This makes it difficult to understand how models arrive at their predictions and makes it challenging for users to trust the model outputs.
@@
LLMs’ tendency to hallucinate can have serious consequences in critical numeric applications. For example, prediction errors in fraud detection in finance can lead to huge financial losses and errors in protection onset of sepsis or cardiac arrest in healthcare can result in patient deaths. Thus, to guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black-box and provide performance guarantees through explaination. Although recent empirical studies Jin et al. (2024); Nie et al. (2023); Liu et al. (2024) demonstrate the benefits of vector representations of embedding learned by LLMs in various numeric downstream tasks, there is little theoretical understanding of their empirical success. Thus, a fundamental question arises: “when (or how) can the next-word prediction capability of LLMs be effectively adapted to numerical domains?"
@@
The main contribution of this paper is to provide an approach to answer this question by exploiting the isotropic property of LLM hidden representations in the contextual embedding space. Isotropy refers to the geometric property wherein vector representations in the embedding space are uniformly distributed in all directions, a characteristic critical for maintaining the expressiveness of the embedding space Arora et al. (2016); Mu and Viswanath (2018). To achieve state-of-the-art performance in numerical domains, we show that the hidden representations of LLMs must exhibit a structured form in contextual embedding space that accounts for the shift-invariance of the softmax function (i.e., the softmax output remains unchanged when all logits are shifted by a constant). Without such structure, the model can shift the logits while keeping the training loss unchanged, thereby leaving the logits ineffective for numeric downstream tasks. By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem of the softmax function.
@@
Our key contributions include: (i) We consider a log-linear model for LLMs and demonstrate theoretically why hidden representations must exhibit structure to address the shift-invariance problem of the softmax function. (ii) We take a deeper look into the hidden representations of pre-trained models and show how isotropy preserves the structural integrity of representations. In particular, we derive an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern and show that $m$ largest eigenvectors of the LLM hidden representations minimize the gradient norm of self-attention. Then by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space. (iii) Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic time series datasets over $6$ different LLMs. Our experiments
@@
demonstrate that different characteristics of numerical data and model architectures could have different impacts on the isotropy.
2 Problem Setup in Numerical Domains
Time Series Tokens and Similarity Measure. Similar to next-word prediction by LLMs, the next-value prediction in the numerical domain can be modeled by time series forecasting techniques Jin et al. (2024); Ansari et al. (2024) which are widely adopted in the machine learning literature. Formally, given a time series $\textbf{x}_{1:T+L}=[x_{1},\ldots,x_{T},\ldots,x_{T+L}]$, where the first $T$ time instances give the historical context, the next $L$ time instances constitute the forecast region, and $x_{t}\in\mathbb{R}$ is the observation of each time instance, we are interested in predicting the joint distribution of next $L$ time instances, $p(\textbf{x}_{T+1:T+L}|\textbf{x}_{1:T})$. Since, the pre-trained models operate on tokens from a finite vocabulary, using them for time series data requires mapping the observations to a finite set of tokens. Based on different numeric applications and LLM architectures, various tokenization techniques, e.g., quantization and scaling Ansari et al. (2024); Rasul et al. (2024), patching Woo et al. (2024); Jin et al. (2024); Nie et al. (2023), and adaptation of language model tokenizer in numeric domians Gruver et al. (2024); Dooley et al. (2023), can be applied to tokenize the time series and create a time series vocabulary $\mathcal{V}$ of $N$ time series tokens, i.e., $|\mathcal{V}|=N$, as shown in Figure 1. Then, the realization of the next $L$ time instances can be obtained by autoregressively sampling from the predicted distribution $p(k_{T+l+1}\,|\,\mathbb{k}_{1:T+l})$, for $l\in\{1,\ldots,L\}$, where $\mathbb{k}_{1:T+l}$ is the tokenized time series and $k_{i}$ be a time series token in time series vocabulary $|\mathcal{V}|$.
Let $\tilde{\mathbb{\Psi}}(k_{i})=\{\mathbb{\psi}_{1}(k_{i}),\mathbb{\psi}_{2}(k_{i}),\ldots\}$ be the set of all LLM contextual embedding instances of time series token $k_{i}$. Here, different contexts in the time series sequences yield different LLM embeddings of $k_{i}$. By constructing $\sum_{k}|\tilde{\mathbb{\Psi}}(k)|=|\mathcal{V}|$, we define the inter-token cosine similarity as,
@@ 2 Problem Setup in Numerical Domains
where $\mathbb{\psi}(k_{i})$ and $\mathbb{\psi}(k_{j})$ are random samples from $\tilde{\mathbb{\Psi}}(k_{i})$. The expectation is taken over all pairs of different tokens. The inter-token cosine similarity metric describes the similarity between different tokens based on the contexts. For ease of reading, we express $T+l$ as $T_{l}$ and $T+l+1$ as $T_{l+1}$ for the rest of the paper.
Model. We consider a general pre-trained model for numerical data and open the black box of the pre-trained model. Specifically, we assume that the observation probability of $k_{T_{l+1}}$ given $\mathbb{k}_{1:T_{l}}$ satisfies the log-linear model Arora et al. (2016)
@@
To define the numeric downstream task, let $z^{*}_{i}(k,l)\!\!:=\!\!\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle$ be the $i$-th logit of the ground-truth model, and assume that the numeric downstream tasks are defined by a function of the logits, i.e., $f^{*}(\mathbb{z^{*}})$. Also let $Z^{*}(k,l)\!\!=\!\!\sum_{i=1}^{N}\exp(z^{*}_{i}(k,l))\!=\!\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)$ be the partition function Arora et al. (2016), i.e., normalization factor. In LLMs, the partition function is often used to normalize the output probabilities of the model, ensuring that they sum to $1$. Then, the normalized ground-truth model $\forall i\!\in\!\mathcal{V}$ is given by
$p(k_{T_{l+1}}\!\!=i\,\,|\,\,\mathbb{k}_{1:T_{l}})$ $=\!\frac{\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}{\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}=\frac{\exp(z^{*}_{i}(k,l))}{Z^{*}(k,l)}$
Since we do not know the ground-truth model in reality, we do not have access to the ground-truth model $\mathbb{\psi}^{*}(k_{i})$ and $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$. Instead, we only have access to the student model $\mathbb{\psi}(k_{i})$ and $\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}})$ that aims to achieve low pre-training loss. We can define the student logits as $\mathbb{z}(k,l)\!\!:=\!\!\{\langle\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}(k_{i})\rangle\}_{i=1}^{|\mathcal{V}|}$. Intuitively, $\mathbb{z}$ are the contextualized representations learned by the student-model during pre-training. Then, the solution of the downstream task is to learn a function $f(k,l)$. Then, the output of the student model $\forall i\!\in\!\mathcal{V}$ can be defines as
@@
Downstream Numerical Task. We consider a simple downstream task whose prediction on categorical distribution is linear in $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$, that is, $f^{*}(k,l)=\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),u^{*}\rangle=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}z^{*}_{i}(k,l),$ where $u^{*}=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}\mathbb{\psi}^{*}(k_{i})\in\mathbb{R}^{D}$ and $a_{j}$ is the coefficient. This model is still not sufficient to provide a performance guarantee to generalize to downstream task in unseen scenarios. However, the log probability difference is proportional to the difference in the value of the perfect model (i.e., ground-truth) $f^{*}(k,l)$. This allows the student model to alter the signs of $f^{*}(k,l)$ without resulting in a large KL divergence Wu et al. (2023). Then, it is more reasonable to model the numeric downstream task as
@@
where $\sigma$ is the ReLU function and $b^{*}_{j}$ denotes the threshold for the logits. The numeric downstream task only considers the logits that are above the threshold, and thus ignores all the entries with very small probabilities.
@@ 3 The Role of Isotropy in Adapting LLMs to Numerical Data
As previously discussed in Section 2, we consider LLM networks whose last layer is usually a softmax layer and the numeric downstream task is determined by the function of the logits. The underlying relation between the logits and softmax function determines the performance of the numeric downstream tasks. However, the softmax function is shift-invariant, that is, the output of the softmax function remains unchanged when all logits are shifted by a constant. Since we do not have any control over the logit shift of the student model on unseen data, good performance during training does not necessarily provide any performance guarantee for the numeric downstream task on unseen scenarios. This can be formalized in the following theorem and the proof is provided in Appendix A.
Theorem 3.1.
Let the logits of the ground-truth model be bounded. Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$. In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numeric downstream tasks.
Theorem 3.1 demonstrates that without any structure in the hidden representations of LLM embeddings, the student model is able to shift the logits for any sample while keeping the pre-training loss unchanged and leaving logits ineffective for the numeric downstream tasks. Consequently, a theoretical guarantee for the numeric downstream task performance needs structure in the LLM representations learned by the pre-trained model. One way to prevent the shift-invariance problem from influencing the performance of the numeric downstream tasks is to keep the partition function stable. Let $\mathbb{\Psi}\!=\!(\mathbb{\psi}_{1}(k),\ldots,\mathbb{\psi}_{|\mathcal{V}|}(k))^{\top}\!\in\!\mathbb{R}^{|\mathcal{V}|\times D}$ be the hidden representations of input time series sequence. Then the stability of the partition function can be assessed through the isotropy in the contextual embedding space Arora et al. (2016); Mu and Viswanath (2018) as follows

### v4

@@ Abstract
Vector representations of contextual embeddings learned by pre-trained large language models (LLMs) are effective in various downstream tasks in numerical domains such as time series forecasting. Despite their significant benefits, the tendency of LLMs to hallucinate in such domains can have severe consequences in applications such as energy, nature, finance, healthcare, retail and transportation, among others. To guarantee prediction reliability and accuracy in numerical domains, it is necessary to open the black box behind the LLM and provide performance guarantees through explanation. However, there is little theoretical understanding of when pre-trained language models help solve numerical downstream tasks. This paper seeks to bridge this gap by understanding when the next-word prediction capability of LLMs can be adapted to numerical domains through a novel analysis based on the concept of isotropy in the contextual embedding space. Specifically, a log-linear model for LLMs is considered in which numerical data can be predicted from its context through a network with softmax in the output layer of LLMs (i.e., language model head in self-attention). For this model, it is demonstrated that, in order to achieve state-of-the-art performance in numerical domains, the hidden representations of the LLM embeddings must possess a structure that accounts for the shift-invariance of the softmax function. By formulating a gradient structure of self-attention in pre-trained models, it is shown how the isotropic property of LLM embeddings in contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem and providing a performance guarantee. Experiments show that different characteristics of numerical data and model architectures have different impacts on isotropy, and this variability directly affects the performances.
@@ 1 Introduction
Large language models (LLMs) have been proven to effective in adapting to various downstream tasks in numerical domains, such as finance Garza and Mergenthaler-Canseco (2023); Yu et al. (2023), energy Gao et al. (2024), climate science Jin et al. (2024), healthcare Wang and Zhang (2024), transportation signals Xu et al. (2024), synthetic tabular generation Dinh et al. (2022); Borisov et al. (2023); Xu et al. (2024), among others. Inspired by the success of pre-trained LLMs, several methods have been developed recently in Gruver et al. (2024); Dooley et al. (2023); Nie et al. (2023); Rasul et al. (2024); Woo et al. (2024); Jin et al. (2024); Ansari et al. (2024) by adapting LLM to numerical domains that deal with time series forecasting. For many of these numerical downstream tasks, training a linear classifier on top of the hidden-layer representations generated by the pre-trained LLMs has been shown to achieve near state-of-the-art performance Jin et al. (2024); Ansari et al. (2024). However, the existing models in Gruver et al. (2024); Dooley et al. (2023); Nie et al. (2023); Rasul et al. (2024); Woo et al. (2024); Jin et al. (2024); Ansari et al. (2024) are treated as a ‘black box’ where numerical forecasts are controlled by complex nonlinear interactions between many parameters. This makes it difficult to understand how models arrive at their predictions and makes it challenging for users to trust the model outputs.
@@
When applied to critical numerical domain use cases, the tendency of LLMs to hallucinate can have serious and detrimental consequences. For example, prediction errors in fraud detection in finance can lead to huge financial losses and errors in protection onset of sepsis or cardiac arrest in healthcare can result in patient deaths. Thus, to guarantee prediction reliability and accuracy in numerical domains, it is necessary to understand the inner working of the so-called black box and provide performance guarantees through explanation. Although recent empirical studies Jin et al. (2024); Nie et al. (2023); Liu et al. (2024) demonstrate the benefits of vector representations of embedding learned by LLMs in various numerical downstream tasks, there is little theoretical understanding of their empirical success. Thus, a fundamental question arises: “When (or how) can the next-word prediction capability of LLMs be effectively adapted to numerical domains?"
@@
The main contribution of this paper is a novel approach for answering this question by exploiting the isotropic property of LLM hidden representations in the contextual embedding space. Isotropy refers to the geometric property wherein vector representations in the embedding space are uniformly distributed in all directions, a characteristic critical for maintaining the expressiveness of the embedding space Arora et al. (2016); Mu and Viswanath (2018). To achieve state-of-the-art performance in numerical domains, we show that the hidden representations of LLMs must exhibit a structured form in contextual embedding space that accounts for the shift-invariance of the softmax function (i.e., the softmax output remains unchanged when all logits are shifted by a constant). Without such structure, the model can shift the logits while keeping the training loss unchanged, thereby leaving the logits ineffective for numerical downstream tasks. By formulating a gradient structure of self-attention in pre-trained models, we show how the isotropic property of LLM embeddings in the contextual embedding space preserves the underlying structure of representations, thereby resolving the shift-invariance problem of the softmax function. In a nutshell, our key contributions include:
  • We consider a log-linear model for LLMs and demonstrate theoretically why hidden representations must exhibit structure to address the shift-invariance problem of the softmax function.
@@
  • We take a deeper look into the hidden representations of pre-trained models and show how isotropy preserves the structural integrity of representations. In particular, we derive an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern and show that the $m$ largest eigenvectors of the LLM hidden representations minimize the gradient norm of self-attention. Then, by projecting the representations into lower dimensions using these $m$ largest eigenvectors, we find the isotropy within the clusters in the contextual embedding space.
  • Finally, we provide a comprehensive evaluation across $12$ real and $10$ synthetic time series datasets over $5$ different LLMs. Our experiments demonstrate that the isotropy of LLM hidden representations varies significantly based on the input data characteristics (i.e., domain, context length and noise level) and model design choices (i.e., tokenization techniques and architecture), which in turn strongly influences forecasting performance in numerical domains.
2 Problem Setup in Numerical Domains
Time Series Tokens and Similarity Measure. Similar to next-word prediction by LLMs, the next-value prediction in the numerical domain can be modeled by time series forecasting techniques which are widely adopted in the machine learning literature Jin et al. (2024); Ansari et al. (2024). Formally, given a time series $\textbf{x}_{1:T+L}=[x_{1},\ldots,x_{T},\ldots,x_{T+L}]$,
@@ 2 Problem Setup in Numerical Domains
where the first $T$ time instances give the historical context, the next $L$ time instances constitute the forecast region, and $x_{t}\in\mathbb{R}$ is the observation of each time instance, we are interested in predicting the joint distribution of the next $L$ time instances, $p(\textbf{x}_{T+1:T+L}|\textbf{x}_{1:T})$. Since, the pre-trained models operate on tokens from a finite vocabulary, using them for time series data requires mapping the observations to a finite set of tokens. Based on different numerical applications and LLM architectures, various tokenization techniques, e.g., quantization and scaling Ansari et al. (2024); Rasul et al. (2024), patching Woo et al. (2024); Jin et al. (2024); Nie et al. (2023), and adaptation of language model tokenizer in numerical domians Gruver et al. (2024); Dooley et al. (2023), can be applied to tokenize the time series and create a time series vocabulary $\mathcal{V}$ of $N$ time series tokens, i.e., $|\mathcal{V}|=N$, as shown in Figure 1. Then, the realization of the next $L$ time instances can be obtained by autoregressively sampling from the predicted distribution $p(k_{T+l+1}\,|\,\mathbb{k}_{1:T+l})$, for $l\in\{1,\ldots,L\}$, where $\mathbb{k}_{1:T+l}$ is the tokenized time series and $k_{i}$ be a time series token in time series vocabulary $|\mathcal{V}|$.
Let $\tilde{\mathbb{\Psi}}(k_{i})=\{\mathbb{\psi}_{1}(k_{i}),\mathbb{\psi}_{2}(k_{i}),\ldots\}$ be the set of all LLM contextual embedding instances of time series token $k_{i}$. Here, different contexts in the time series sequences yield different LLM embeddings of $k_{i}$. By constructing $\sum_{k}|\tilde{\mathbb{\Psi}}(k)|=|\mathcal{V}|$, we define the inter-token cosine similarity as:
@@
where $\mathbb{\psi}(k_{i})$ and $\mathbb{\psi}(k_{j})$ are random samples from $\tilde{\mathbb{\Psi}}(k_{i})$. The expectation is taken over all pairs of different tokens. The inter-token cosine similarity metric describes the similarity between different tokens based on the contexts. For notational simplicity, we express $T+l$ as $T_{l}$ and $T+l+1$ as $T_{l+1}$ hereinafter.
Model. We consider a general pre-trained model for numerical data and open the black box of the pre-trained model. Specifically, we assume that the observation probability of $k_{T_{l+1}}$ given $\mathbb{k}_{1:T_{l}}$ satisfies the log-linear model as in Arora et al. (2016)
@@
To define the numerical downstream task, let $z^{*}_{i}(k,l)\!\!:=\!\!\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle$ be the $i$-th logit of the ground-truth model, and assume that the numerical downstream tasks are defined by a function of the logits, i.e., $f^{*}(\mathbb{z^{*}})$. Also let $Z^{*}(k,l)\!\!=\!\!\sum_{i=1}^{N}\exp(z^{*}_{i}(k,l))\!=\!\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)$ be the partition function Arora et al. (2016), i.e., normalization factor. In LLMs, the partition function is often used to normalize the output probabilities of the model, ensuring that they sum to $1$. Then, the normalized ground-truth model $\forall i\!\in\!\mathcal{V}$ is given by
$p(k_{T_{l+1}}\!\!=i\,\,|\,\,\mathbb{k}_{1:T_{l}})$ $=\!\frac{\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}{\sum_{i=1}^{|\mathcal{V}|}\exp(\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}^{*}(k_{i})\rangle)}=\frac{\exp(z^{*}_{i}(k,l))}{Z^{*}(k,l)}.$
Since we do not know the ground-truth model in reality, we do not have access to the ground-truth model components $\mathbb{\psi}^{*}(k_{i})$ and $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$. Instead, we only have access to the student model $\mathbb{\psi}(k_{i})$ and $\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}})$ that aims to achieve low pre-training loss. We can define the student logits as $\mathbb{z}(k,l)\!\!:=\!\!\{\langle\mathbb{\mathbb{\psi}}(\mathbb{k}_{1:T_{l}}),\mathbb{\psi}(k_{i})\rangle\}_{i=1}^{|\mathcal{V}|}$. Intuitively, $\mathbb{z}$ are the contextualized representations learned by the student-model during pre-training. Then, the solution of the downstream task is to learn a function $f(k,l)$. Then, the output of the student model $\forall i\!\in\!\mathcal{V}$ can be defines as
@@
Downstream Numerical Task. We consider a simple downstream task whose prediction on categorical distribution is linear in $\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}})$, that is, $f^{*}(k,l)=\langle\mathbb{\mathbb{\psi}}^{*}(\mathbb{k}_{1:T_{l}}),u^{*}\rangle=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}z^{*}_{i}(k,l),$ where $u^{*}=\sum_{i=1}^{|\mathcal{V}|}a^{*}_{i}\mathbb{\psi}^{*}(k_{i})\in\mathbb{R}^{D}$ and $a_{j}$ is the coefficient. This model is still not sufficient to provide a performance guarantee to generalize to downstream task in unseen scenarios. However, the log probability difference is proportional to the difference in the value of the perfect model (i.e., ground-truth) $f^{*}(k,l)$. This allows the student model to alter the signs of $f^{*}(k,l)$ without resulting in a large KL divergence Wu et al. (2023). Then, it is more reasonable to model the numerical downstream task as
@@
where $\sigma$ is the ReLU function and $b^{*}_{j}$ denotes the threshold for the logits. The numerical downstream task only considers the logits that are above the threshold, and thus ignores all the entries with very small probabilities.
Despite the empirical success of language models in numerical domains, there remains a fundamental gap in understanding when and why these models generalize reliably to numeric downstream tasks such as time series forecasting across different numerical domains. A key challenge lies in the mismatch between training and inference behavior, i.e., good training performance does not always guarantee robust performance at inference time. To address this challenge, we propose a novel theoretical framework grounded in the isotropic property of the contextual embedding space. We show that the presence of strong isotropy in LLM hidden representations stabilizes the partition function, effectively resolving the softmax shift-invariance problem and leading to reliable inference performance. The next section formalizes this insight and provides theoretical justification for using isotropy as a key indicator of model reliability in numerical settings.
@@ 3 The Role of Isotropy in Adapting LLMs to Numerical Data
As discussed in Section 2, we consider LLM networks whose last layer is usually a softmax layer and the numerical downstream task is determined by the function of the logits. The underlying relation between the logits and softmax function determines the performance of the numerical downstream tasks. However, the softmax function is shift-invariant, that is, the output of the softmax function remains unchanged when all logits are shifted by a constant. Since we do not have any control over the logit shift of the student model on unseen data, good performance during training does not necessarily provide any performance guarantee for the numerical downstream task on unseen scenarios. The lack of perfromance guarantee under uncontrolled logit shifts on unseed data can be formalized in the following theorem:
Theorem 1.
Let the logits of the ground-truth model be bounded. Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$. In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numerical downstream tasks.
Proof.
The proof is provided in Appendix A. ∎
Theorem 1 shows that without any structure in the hidden representations of LLM embeddings, student model can shift the logits for any sample while keeping the pre-training loss unchanged and leaving logits ineffective for the numerical downstream tasks. Consequently, a theoretical guarantee for numerical downstream task performance will require structure in the LLM representations learned by the pre-trained model.
In this paper, we make an observation that to prevent the shift-invariance problem from influencing the performance of the numerical downstream tasks, it is necessary to keep the partition function stable. Let $\mathbb{\Psi}\!=\!(\mathbb{\psi}_{1}(k),\ldots,\mathbb{\psi}_{|\mathcal{V}|}(k))^{\top}\!\in\!\mathbb{R}^{|\mathcal{V}|\times D}$ be the hidden representations of input time series sequence. Then, the stability of the partition function can be assessed through the isotropy in the contextual embedding space Arora et al. (2016); Mu and Viswanath (2018) as follows
@@
Building on this theoretical foundation, we now turn to the following empirical question: “How can we measure and interpret isotropy in practice, and how does it relate to generalization across numerical domains?". Motivated by Theorem 3.1 and the need for structural constraints in LLM representations, we analyze the effective dimensionality and cluster organization of LLM’s hidden representations in the contextual embedding space. These analyses reveal how isotropy manifests in pre-trained LLMs and how its presence correlates with the model’s ability to generalize to time series forecasting across different numerical domains, and hence, provides a performance guarantee. Section 4 introduces methods for quantifying this internal structural integrity using spectral alignment mechanism, principal component analysis (PCA) and cluster-based isotropy metric, and consequently, linking theoretical reliability (i.e., performance guarantee) to empirical generalizability.